OpenStudy (anonymous):

two balls fall freely from the heights in the ratio 1:4 Find the ratio of times for the ball to reach the ground.

OpenStudy (anonymous):

\[s= ut +\frac{ 1 }{ 2 } at ^{2} \] Use this kinematic relation. s=displacement u=initial velocity a=acceleration=g in this case t=time For you question u=0 hence \[s= \frac{ 1 }{ 2 } a t ^{2}\] or \[\frac{ s1 }{ s2 } = (\frac{ t1 }{ t2 })^{2} \] s1/s2 is given in the question=1/4 You can find t1/t2 like this = 1/2